# MHR: Momentum Human Rig

### v1

@@ 1 Introduction
The landscape of digital human modeling has rapidly evolved, driving progress across fields such as virtual reality avatar creation lombardi2019neural; saito2020pifuhd; weng2022humannerf; xiu2022icon, advanced motion capture yin2023hi4d; cheng2023dna; peng2021neural; zheng2022structured, realistic simulation of human-object interactions bhatnagar2022behave; chao2015hico; wang2024review, and generative character synthesis peng2024charactergen; ren2023make; liang2024rich. Central to these innovations are parametric body models SMPL:2015; SMPL-X:2019; xu2020ghum; anguelov2005scape; wang2020blsm; yang2020facescape, which translate compact shape and pose descriptors into detailed, articulated meshes. The ability to flexibly and accurately represent human form and movement is essential for enabling new applications and deepening our understanding of human-centric data.
@@ 7 Author Contributions
Kris Kitani - SAM3D integration, ATLAS development

### v2

@@ 1 Introduction
The landscape of digital human modeling has rapidly evolved, driving progress across fields such as avatar creation lombardi2019neural; saito2020pifuhd; weng2022humannerf; xiu2022icon, motion capture yin2023hi4d; cheng2023dna; peng2021neural; zheng2022structured, simulation of human-object interactions bhatnagar2022behave; chao2015hico; wang2024review, and generative character synthesis peng2024charactergen; ren2023make; liang2024rich. Central to these innovations are parametric body models SMPL:2015; SMPL-X:2019; xu2020ghum; anguelov2005scape; wang2020blsm; yang2020facescape, which translate compact shape and pose descriptors into articulated meshes. The ability to flexibly and accurately represent human form and movement is essential for enabling new applications and deepening our understanding of human-centric data.
@@ 7 Author Contributions
Kris Kitani - Open sourcing, SAM3D integration, ATLAS development